OpenStudy (anonymous):

|-13| + 5^2 = 80y - |-37| - 79y

OpenStudy (bahrom7893):

13+25 = 80y - 37 - 79y 13+25 + 37=y

OpenStudy (bahrom7893):

y = 38+37= 75